Circular No. 6667

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997cn IN NGC 5490 M. Turatto and S. Benetti, European Southern Observatory (ESO);

and C. Pereira and L. Da Silva, Observatorio Nacional, Rio de Janeiro, report: “Inspection of a fully-reduced CCD spectrogram (range 354–741 nm, resolution 0.6 nm) obtained on May 22.13 UT with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph) at La Silla confirms that SN 1997cn is a peculiar type-Ia supernova. The spectrum closely resembles that of SN 1991bg one day after maximum (cf. Turatto et al. 1991, MNRAS 283, 1). In particular, the spectrum shows the broad absorption between 420 and 450 nm that is attributed to Ti II. The unusual profiles of S II 546.8-, 561.2-, and 565.4-nm lines are also closely matched. The expansion velocity deduced from the minimum of the Si II 635.5-nm line is low (v = 9400 km/s) in comparison to that of SN 1991bg at the same phase. From the spectrophotometry, we estimate V ∼ 16.7 and B − V ∼ +0.9, corresponding to M_{v}(max) ∼ −17.4 (Hubble constant 75 km s^{−1} Mpc^{−1}). Note that the parent galaxy is an elliptical (as in the case of SN 1991bg), and there is no evident signature of Na I D absorption. Therefore, the faint absolute magnitude and the red color are intrinsic.”

COMETS 6P/D’ARREST AND C/1995 Q1 (BRADFIELD) M. Abbott, Center for Extreme Ultraviolet Astrophysics, Berkeley;

and V. A. Krasnopolsky and M. J. Mumma, Goddard Space Flight Center, NASA, report: “We searched for soft-x-ray radiation from comet 6P on 1995 Sept. 4–5 (r = 1.42 AU, Δ = 0.46 AU) and from comet C/1995 Q1 on 1995 Nov. 6–7 (r = 1.50 AU, Δ = 1.26 AU). The observations were made in the range 70–180 eV with the Extreme Ultraviolet Explorer orbiting observatory. We detected comet 6P at a luminosity equal to 1.1 × 10^{24} photons/s within an aperture radius of 120 000 km. No emission was detected from C/1995 Q1, with an upper limit of 3 × 10^{23} photons/s in the same aperture.”

COMET C/1995 O1 (HALE-BOPP) Total visual magnitude estimates (naked eye unless otherwise noted):

Apr. 23.82, −0.2 (M. V. Zanotta, Milan, Italy); May 1.87, 0.1 (R. J. Bouma, Groningen, The Netherlands); 6.83, 0.4 (Zanotta); 9.34, 0.4 (T. Lovejoy, Brisbane, Queensland); 16.08, 1.2 (G. W. Kronk, Troy, IL); 20.09, 1.6 (Kronk, 7×35 binoculars); 22.09, 2.1: (Kronk, 20×80 binoculars).

1997 May 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green